3. Suppose that the demand for widgets has the following elasticities:
Older person’s price elasticity of demand = 4.0
Price elasticity of demand by all others = 2.0
Market price elasticity of demand = 3.0
All demanders are either older people, or others.

10 a. If the MC = AC = 4 for widgets, what price would a simple monopolist charge?
20 b. Suppose that the monopolist attempts to price discriminate between the two groups of demanders. What problems would the monopolist have in successfully charging different prices to them?
10 c. How might it solve one of those problems?
1
Expert's answer
2014-05-14T11:10:57-0400
10. If the MC = AC = 4 for widgets, Ed = 3, themonopolist’s pricing rule as a function of the elasticity of demand for its
product is: (P - MC)/P = 1/Ed
1 - 4/P = 1/3
4/P = 2/3
P = $6

20. There will be a problem how to charge different prices and how to dispose of the reselling widgets between different price groups.

10. The monopolist may create different markets for older an the others and to make recelling widgets to another group impossible, making these markets unavailable to younger or older persons, requiring the selling of widgets to persons with appropriate age.